# Multimodal Instruction Tuning with Conditional Mixture of LoRA

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable proficiency in diverse tasks across different domains, with an increasing focus on improving their zero-shot generalization capabilities for unseen multimodal tasks. Multimodal instruction tuning has emerged as a successful strategy for achieving zero-shot generalization by fine-tuning pre-trained models on diverse multimodal tasks through instructions. As MLLMs grow in complexity and size, the need for parameter-efficient fine-tuning methods like Low-Rank Adaption (LoRA), which fine-tunes with a minimal set of parameters, becomes essential. However, applying LoRA in multimodal instruction tuning presents the challenge of task interference, which leads to performance degradation, especially when dealing with a broad array of multimodal tasks. To address this, this paper introduces a novel approach that integrates multimodal instruction tuning with Conditional Mixture-of-LoRA (MixLoRA). It innovates upon LoRA by dynamically constructing low-rank adaptation matrices tailored to the unique demands of each input instance, aiming to mitigate task interference. Experimental results on various multimodal evaluation datasets indicate that MixLoRA not only outperforms the conventional LoRA with the same or even higher ranks, demonstrating its efficacy and adaptability in diverse multimodal tasks.
@@
Ying ${}^{\spadesuit}$  Zhiyang ${}^{\spadesuit}$ Qifan ${}^{\diamondsuit}$u ${}^{\clubsuit}$enpeng ${}^{\heartsuit}$ifu ${}^{\spadesuit}$$^{\spadesuit}$irginia Tech  ${}^{\diamondsuit}$eta AI  ${}^{\clubsuit}$ice University  ${}^{\heartsuit}$he Pennsylvania State University {yings,lifuh}@vt.edu
@@ 2 Related Work
Instruction tuning Wei et al. (2021) significantly improves the generalization of large language models to unseen tasks based on natural language instructions. With the advent of multimodal large language models, the scope of instruction tuning has expanded to encompass multimodal and vision tasks, facilitated by the development of diverse multimodal instruction datasets, including both machine-generated Liu et al. (2023); Zhao et al. (2023); Zhu et al. (2023); Yin et al. (2023); Li et al. (2023b); Ye et al. (2023) and human-annotated Xu et al. (2023b). Recently, Vision-Flan Xu et al. (2023a) stands out as a comprehensive human-annotated visual instruction tuning dataset, covering a wide range of 187 tasks, making it ideal for our training.
@@
Parameter-efficient fine-tuning (PEFT) Lee et al. (2019); Hu et al. (2021); Li and Liang (2021); Karimi Mahabadi et al. (2021); Guo et al. (2021); Zaken et al. (2022) strategies have become key in efficiently adapting large pre-trained models to various downstream tasks with minimal parameter adjustments. Among these, LoRA Hu et al. (2021) demonstrates competitive trade-offs between performance and parameter efficiency, making it widely adopted. PEFT methods typically utilize shared adaptation parameters across diverse tasks or train task-specific adapters. However, when applied to multimodal instruction tuning, which requires simultaneous adaptation to diverse instruction tasks, PEFT can encounter task interference, highlighting the need for more adaptable and versatile PEFT methods to adeptly handle the complexities of multimodal instruction tuning.
@@ 3.1 Background: Low-Rank Adaptation
Low-Rank Adaptation (LoRA) Hu et al. (2021) is a parameter-efficient fine-tuning method that fine-tunes only the trainable rank decomposition matrices injected in each layer of the Transformer Vaswani et al. (2017). As illustrated in Figure 1 (a), consider a linear layer, represented by $\tilde{h}=Wh$, where $W\in\mathbb{R}^{d_{out}\times d_{in}}$ denotes the pre-trained weight, with $d_{in}$ and $d_{out}$ being the input and output dimensions, respectively. LoRA modifies the model parameters by injecting low-rank decomposition matrices as the weight adjustment matrices, which can be expressed as: {gather+}[0.8] ~h = W h + ΔW h = W h + α⋅BA h, where $\Delta W=BA$ represents the trainable weight adjustment matrices formed by low-rank matrices $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$, with the rank $r\ll\min(d_{in},d_{out})$. The scalar $\alpha\geq 1$ controls the influence of the weight adjustment matrices. During fine-tuning, only these low-rank decomposition matrices, referred to as LoRA $A$ and LoRA $B$ throughout this paper, are updated, allowing for rapid, task-specific adaptation by training distinct LoRA $A$ and $B$ for each downstream task.
@@ 3.2 Investigating Task Interference in Multimodal Instruction Tuning
Our study delves into task interference in parameter-efficient multimodal instruction tuning by analyzing gradient direction conflicts between task pairs. For each task pair $i$ and $j$, we first estimate the change in loss $L_{i}$ of task $i$, when optimizing the shared parameters $\theta$ according to the loss $L_{j}$ of task $j$, following Zhu et al. (2022): {gather+}[0.8] Δ_j L_i(x_i) = E_x_j ( L_i(x_i; θ) - L_i(x_i; θ- λθLj(xj)∥ ∇θLj(xj) ∥ ) )
λE_x_j ( ∇θLj(xj)∥ ∇θLj(xj) ∥^T ∇_θ L_i(x_i) ) where $x_{i}$ and $x_{j}$ are sampled training batches for tasks $i$ and $j$, and $\lambda$ is the learning rate.
The interference of task $j$ on task $i$ is then quantified as follows: {gather+}[0.8] I_i, j = E_x_i (ΔjLi(xiiLi(xi))
@@ 4 Conditional Mixture-of-LoRA
The core of Conditional Mixture-of-LoRA lies in representing the weight adjustment matrices $\Delta W$ from Equation 3.1 via tensor decomposition: {gather+}[0.8] ΔW = BA = ∑_i=1^r b_i ⊗a_i, where $\{a_{i},b_{i}\}_{i=1}^{r},a_{i}\in\mathbb{R}^{d_{in}\times 1},b_{i}\in\mathbb{R}^{d_{out}\times 1}$ are the rank $r$ decomposition factors of $\Delta W$.
Leveraging the concept that $\Delta W$ can be expressed as sum of outer products of low-rank decomposition factors $a_{i}$ and $b_{i}$, MixLoRA introduces a Dynamic Factor Selection module. This module dynamically constructs unique $\Delta W$ for specific inputs by selecting $r$ appropriate factors from an expanded pool of decomposition factors $\{a_{e}\}_{e=1}^{E},\{b_{e}\}_{e=1}^{E},E>r$, as shown in Fig. 1 (b).
@@ 4.1.1 Independent Factor Selection
MixLoRA employs two Independent Factor Selection (IFS) routers, $R^{A}_{\text{IFS}}(\cdot)$ and $R^{B}_{\text{IFS}}(\cdot)$, to select $r$ relevant factors for LoRA $A$ and $B$, respectively, as shown in Figure 3. IFS routers employ an instance-based routing method, which is more memory-efficient than conventional input-token-based routing, for selecting $r$ decomposition factors. The routing strategy can be expressed as: {gather+}[0.8] R^A_IFS(h) = Avg(h), where $\text{Avg}(\cdot)$ averages across the sequence dimension of the hidden states $h\in\mathbb{R}^{seq\times d_{in}}$ from the preceding layer.
@@
The factor selection process involves calculating vectors $g_{A}\in\mathbb{R}^{E}$ and $g_{B}\in\mathbb{R}^{E}$ to selectively identify specific subsets of decomposition factors from the set $\{a_{e}\}_{e=1}^{E}$ and $\{b_{e}\}_{e=1}^{E}$, respectively. To compute $g_{A}$, input $R^{A}_{\text{IFS}}(h)\in\mathbb{R}^{d_{in}}$ is processed through a dense layer with weights $W_{A}\in\mathbb{R}^{E\times d_{in}}$, followed by a softmax normalization and top-$k$ selection: {gather+}[0.8] g_A = top_r(softmax(W_A ⋅R^A_IFS(h))).
@@ 4.1.2 Conditional Factor Selection
With the IFS router, LoRA $A$ is assembled from chosen decomposition factors, denoted as $A=[a_{1},\cdots,a_{r}]^{T}$, where $A\in\mathbb{R}^{r\times d_{in}}$. Following this, the CFS router employs a weight tensor $W_{AB}\in\mathbb{R}^{r\times d_{in}\times E}$ to map each factor $A[i]\in\mathbb{R}^{1\times d_{in}}$ in $A$ to an expert dimension $E$. The mapping process for each factor $A[i]$, normalized via softmax and aggregated across $r$ factors, is given by: {gather+}[0.8] R^B_CFS(A) = ∑_i=1^r softmax(A[i] ⋅W_AB[i]), where $W_{AB}[i]\in\mathbb{R}^{d_{\text{in}}\times E}$ is the mapping matrix associated with $A[i]$.
The factors selection for LoRA $B$ integrates outputs from both the IFS $R^{B}_{\text{IFS}}(\cdot)$ and CFS $R^{B}_{\text{CFS}}(\cdot)$ routers via a late fusion strategy, forming the selection vector $g_{B}$ as follows: {gather+}[0.8] p^B_IFS = softmax(W^B_IFS ⋅R^B_IFS(h))
p^B_CFS = softmax(R^B_CFS(A))
g_B = top_r(p^B_IFS + p^B_CFS).
The final selection vector $g_{B}$ is determined by combining the probability distributions $p^{B}_{\text{IFS}}$ and $p^{B}_{\text{CFS}}$ from the IFS and CFS routers. This CFS strategy enables the selection for LoRA $B$ to be informed by factors selected for LoRA $A$, fostering a more cohesive selection process.
@@ 4.1.3 Reconstruction of Dynamic Adaptation Matrices
Finally, MixLoRA constructs dynamic adaptation matrices by leveraging the factor selection vectors $g_{A}$ and $g_{B}$, gathering the chosen factors ${a_{k},b_{k}}{k\in K},|K|=r$, to assemble the final matrices for LoRA $A$ and $B$. Consequently, in each forward pass, the weight adjustment matrix $\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$ is dynamically calculated based on these selected factors, formulated as: {gather+}[0.8] ΔW = BA = [b_1,⋯, b_r] [a_1,⋯, a_r]^T
@@ 5.1 Datasets
We perform instruction tuning on Vision-Flan Xu et al. (2023a), a human-annotated multimodal instruction tuning dataset with 187 diverse tasks. Its diversity in visual instruction tasks makes it ideal for investigating task interference. To minimize computational cost, we utilize a scaled-down version with up to 1,000 instances per task, totaling 191,105 instances.
@@
We evaluate our method on MME Fu et al. (2023), a comprehensive multi-modal evaluation benchmark measuring both perception and cognition abilities across 14 subtasks. Alongside MME, we further probe the model’s various capabilities using 7 multimodal datasets. For Optical Character Recognition, we utilize Text-VQA Singh et al. (2019), and for reasoning, we employ the Visual Spatial Reasoning (VSR) dataset Liu et al. (2022). Perception capability is tested on CIFAR-10/100 Krizhevsky et al. (2009) and MNIST LeCun (1998), following the guidance of Zhai et al. (2023). The SNLI-VE dataset Xie et al. (2019) evaluates the Visual Entailment capabilities, and POPE Li et al. (2023c) examines the tendency to objects hallucination.
@@ 6 Results and Discussion
Moreover, we visualize the task interference scores using Equation 3.2 and 3.2. Given that MixLoRA dynamically selects a subset of factors ($r$ out of $E$) for different instances, we record gradients concerning all $E$ factors and compare the task interference scores between standard LoRA models (with $r$=16) and MixLoRA (with $E$ = 16 and $r$ = 4). Figure 7 visualizes the interference scores for both LoRA $A$ and LoRA $B$ aggregated across all adaptation layers, including MLP and self-attention layers. The analysis reveals that MixLoRA ($E$=16, $r$=4) exhibits lower negative interference scores compared to the standard LoRA ($r$=16), underscoring MixLoRA’s efficacy in reducing task interference.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable proficiency in diverse tasks across different domains, with an increasing focus on improving their zero-shot generalization capabilities for unseen multimodal tasks. Multimodal instruction tuning has emerged as a successful strategy for achieving zero-shot generalization by fine-tuning pre-trained models on diverse multimodal tasks through instructions. As MLLMs grow in complexity and size, the need for parameter-efficient fine-tuning methods like Low-Rank Adaption (LoRA), which fine-tunes with a minimal set of parameters, becomes essential. However, applying LoRA in multimodal instruction tuning presents the challenge of task interference, which leads to performance degradation, especially when dealing with a broad array of multimodal tasks. To address this, this paper introduces a novel approach that integrates multimodal instruction tuning with Conditional Mixture-of-LoRA (MixLoRA). It innovates upon LoRA by dynamically constructing low-rank adaptation matrices tailored to the unique demands of each input instance, aiming to mitigate task interference. Experimental results on various multimodal evaluation datasets indicate that MixLoRA not only outperforms the conventional LoRA with the same or even higher ranks, demonstrating its efficacy and adaptability in diverse multimodal tasks [^1] Our code is publicly available at https://github.com/VT-NLP/MixLoRA.
@@
Ying Shen  Zhiyang Xu Qifan WangYu ChengWenpeng YinLifu Huang Virginia Tech  Meta AI  The Chinese University of Hong Kong The Pennsylvania State University {yings, zhiyangx, lifuh}@vt.eduwqfcr@meta.com chengyu@cse.cuhk.edu.hkwfy5054@psu.edu
@@ 2 Related Work
Instruction tuning Wei et al. (2021) significantly improves the generalization of large language models to unseen tasks based on natural language instructions Ouyang et al. (2022); Taori et al. (2023). With the advent of multimodal large language models, the scope of instruction tuning has expanded to encompass multimodal and vision tasks, facilitated by the development of diverse multimodal instruction datasets, including both machine-generated Liu et al. (2023); Zhao et al. (2023); Zhu et al. (2023); Yin et al. (2023); Li et al. (2023b); Ye et al. (2023) and human-annotated Xu et al. (2023b). Recently, Vision-Flan Xu et al. (2023a) stands out as a comprehensive human-annotated visual instruction tuning dataset, covering a wide range of 187 tasks, making it ideal for our training.
@@
Parameter-efficient fine-tuning (PEFT) Lee et al. (2019); Hu et al. (2021); Li and Liang (2021); Karimi Mahabadi et al. (2021); Guo et al. (2021); Zaken et al. (2022) strategies have become key in efficiently adapting large pre-trained models to various downstream tasks with minimal parameter adjustments. These methods typically involve freezing most of the pre-trained model while fine-tuning a small fraction of the parameters to facilitate the adaptation process. Among these, LoRA Hu et al. (2021) demonstrates competitive trade-offs between performance and parameter efficiency, making it widely adopted. PEFT methods typically utilize shared adaptation parameters across diverse tasks or train task-specific adapters. However, when applied to multimodal instruction tuning, which requires simultaneous adaptation to diverse instruction tasks, PEFT can encounter task interference, highlighting the need for more adaptable and versatile PEFT methods to adeptly handle the complexities of multimodal instruction tuning.
@@ 3.1 Background: Low-Rank Adaptation
Low-Rank Adaptation (LoRA) Hu et al. (2021) is a parameter-efficient fine-tuning method that fine-tunes only the trainable rank decomposition matrices injected in each layer of the Transformer Vaswani et al. (2017). As illustrated in Figure 1 (a), consider a linear layer, represented by $\tilde{h}=Wh$, where $W\in\mathbb{R}^{d_{out}\times d_{in}}$ denotes the pre-trained weight, with $d_{in}$ and $d_{out}$ being the input and output dimensions, respectively. LoRA modifies the model parameters by injecting low-rank decomposition matrices as the weight adjustment matrices, which can be expressed as:
$\tilde{h}=Wh+\Delta Wh=Wh+\alpha\cdot BAh,$ (2)
where $\Delta W=BA$ represents the trainable weight adjustment matrices formed by low-rank matrices $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$, with the rank $r\ll\min(d_{in},d_{out})$. The scalar $\alpha\geq 1$ controls the influence of the weight adjustment matrices. During fine-tuning, only these low-rank decomposition matrices, referred to as LoRA $A$ and LoRA $B$ throughout this paper, are updated, allowing for rapid, task-specific adaptation by training distinct LoRA $A$ and $B$ for each downstream task.
@@ 3.2 Investigating Task Interference in Multimodal Instruction Tuning
Our study delves into task interference in parameter-efficient multimodal instruction tuning by analyzing gradient direction conflicts between task pairs. For each task pair $i$ and $j$, we first estimate the change in loss $L_{i}$ of task $i$, when optimizing the shared parameters $\theta$ according to the loss $L_{j}$ of task $j$, following Zhu et al. (2022): {gather+}[0.84] Δ_j L_i(x_i) = E_x_j ( L_i(x_i; θ) - L_i(x_i; θ- λθLj(xj)∥ ∇θLj(xj) ∥ ) )
λE_x_j ( ∇θLj(xj)∥ ∇θLj(xj) ∥^T ∇_θ L_i(x_i) ), where $x_{i}$ and $x_{j}$ are sampled training batches for tasks $i$ and $j$, and $\lambda$ is the learning rate.
The interference of task $j$ on task $i$ is then quantified as follows:
$\mathcal{I}_{i,j}=\mathbb{E}_{x_{i}}\left(\frac{\Delta_{j}L_{i}(x_{i})}{\Delta_{i}L_{i}(x_{i})}\right).$ (3)
@@ 4 Conditional Mixture-of-LoRA
The core of Conditional Mixture-of-LoRA lies in representing the weight adjustment matrices $\Delta W$ from Equation 2 via tensor decomposition:
$\Delta W=BA=\sum_{i=1}^{r}b_{i}\otimes a_{i},$ (4)
where $\otimes$ denotes the outer product and $\{a_{i},b_{i}\}_{i=1}^{r},a_{i}\in\mathbb{R}^{d_{in}\times 1},b_{i}\in\mathbb{R}^{d_{out}\times 1}$ are the rank $r$ decomposition factors of $\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$.
Leveraging the concept that $\Delta W$ can be expressed as the sum of outer products of low-rank decomposition factors $a_{i}$ and $b_{i}$, MixLoRA introduces a Dynamic Factor Selection module. This module dynamically constructs unique $\Delta W$ for specific inputs by selecting $r$ appropriate factors from an expanded pool of decomposition factors $\{a_{e}\}_{e=1}^{E},\{b_{e}\}_{e=1}^{E},E>r$, as shown in Fig. 1 (b). Here, the number of factors $E$ is much larger than the rank $r$. Following LoRA Hu et al. (2021), we use a random Gaussian initialization for $\{a_{e}\}_{e=1}^{E}$ and zero for $\{b_{e}\}_{e=1}^{E}$.
@@ 4.1.1 Independent Factor Selection
MixLoRA employs two Independent Factor Selection (IFS) routers, $R^{A}_{\text{IFS}}(\cdot)$ and $R^{B}_{\text{IFS}}(\cdot)$, to select $r$ relevant factors for LoRA $A$ and $B$, respectively, as shown in Figure 3. IFS routers employ an instance-based routing method, which is more memory-efficient than conventional input-token-based routing, for selecting $r$ decomposition factors. The routing strategy can be expressed as:
$R^{A}_{\text{IFS}}(h)=\text{Avg}(h),$ (5)
where $\text{Avg}(\cdot)$ averages across the sequence dimension of the hidden states $h\in\mathbb{R}^{seq\times d_{in}}$ from the preceding layer, resulting $R^{A}_{\text{IFS}}(h)\in\mathbb{R}^{d_{in}}$.
@@
The factor selection process involves calculating vectors $g_{A}\in\mathbb{R}^{E}$ and $g_{B}\in\mathbb{R}^{E}$ to selectively identify specific subsets of decomposition factors from the set $\{a_{e}\}_{e=1}^{E}$ and $\{b_{e}\}_{e=1}^{E}$, respectively. To compute $g_{A}$, input $R^{A}_{\text{IFS}}(h)\in\mathbb{R}^{d_{in}}$ is processed through a dense layer with weights $W_{A}\in\mathbb{R}^{E\times d_{in}}$, followed by a softmax normalization and top-$k$ selection:
$g_{A}=\text{top}_{r}(\text{softmax}(W_{A}\cdot R^{A}_{\text{IFS}}(h))).$ (6)
@@ 4.1.2 Conditional Factor Selection
With the IFS router, LoRA $A$ is assembled from chosen decomposition factors, denoted as $A=[a_{1},\cdots,a_{r}]^{T}$, where $A\in\mathbb{R}^{r\times d_{in}}$. Following this, the CFS router employs a weight tensor $W_{AB}\in\mathbb{R}^{r\times d_{in}\times E}$ to map each factor $A[i]\in\mathbb{R}^{1\times d_{in}}$ in $A$ to an expert dimension $E$. The mapping process for each factor $A[i]$, normalized via softmax and aggregated across $r$ factors, is given by:
$R^{B}_{\text{CFS}}(A)=\sum_{i=1}^{r}\text{softmax}(A[i]\cdot W_{AB}[i]),$ (7)
where $W_{AB}[i]\in\mathbb{R}^{d_{\text{in}}\times E}$ is the mapping matrix associated with $A[i]$.
The factors selection for LoRA $B$ integrates outputs from both the IFS $R^{B}_{\text{IFS}}(\cdot)$ and CFS $R^{B}_{\text{CFS}}(\cdot)$ routers via a late fusion strategy, forming the selection vector $g_{B}$ as follows:
$p^{B}_{\text{IFS}}=\text{softmax}(W^{B}_{\text{IFS}}\cdot R^{B}_{\text{IFS}}(h))$ (8)
$p^{B}_{\text{CFS}}=\text{softmax}(R^{B}_{\text{CFS}}(A))$ (9)
$g_{B}=\text{top}_{r}(p^{B}_{\text{IFS}}+p^{B}_{\text{CFS}}),$ (10)
where, $R^{B}_{\text{IFS}}(h)\in\mathbb{R}^{d_{out}}$ for LoRA $B$ is computed similarly to $R^{A}_{\text{IFS}}(h)$.
The probability distribution $p^{B}_{\text{IFS}}\in\mathbb{R}^{E}$ and $p^{B}_{\text{CFS}}\in\mathbb{R}^{E}$ reflect the selections from the independent and conditional routers, respectively. The final selection vector $g_{B}\in\mathbb{R}^{E}$ is determined by combining these distributions and identifying the top $r$ factors. This CFS strategy enables the selection for LoRA $B$ to be informed by factors selected for LoRA $A$, fostering a more cohesive selection process.
@@ 4.1.3 Reconstruction of Dynamic Adaptation Matrices
Finally, MixLoRA constructs dynamic adaptation matrices by leveraging the factor selection vectors $g_{A}$ and $g_{B}$, gathering the chosen factors $\{a_{k},b_{k}\}_{k\in K},|K|=r$, to assemble the final matrices for LoRA $A$ and $B$. Consequently, in each forward pass, the weight adjustment matrix $\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$ is dynamically calculated based on these selected factors, formulated as:
$\Delta W=BA=[b_{1},\cdots,b_{r}][a_{1},\cdots,a_{r}]^{T}.$ (11)
@@ 5.1 Datasets
To validate the effectiveness of MixLoRA, we perform instruction tuning on Vision-Flan Xu et al. (2023a), a human-annotated multimodal instruction tuning dataset with 187 diverse tasks. Its diversity in visual instruction tasks makes it ideal for investigating task interference. To minimize computational cost, we utilize a scaled-down version with up to 1,000 instances per task, totaling 182,167 instances.
@@
We evaluate our method on MME Fu et al. (2023), a comprehensive multi-modal evaluation benchmark measuring both perception and cognition abilities across 14 subtasks, to assess the overall capabilities of MixLoRA. Alongside MME, we further probe the model’s various capabilities using 7 multimodal datasets. For Optical Character Recognition, we utilize Text-VQA Singh et al. (2019), and for reasoning, we employ the Visual Spatial Reasoning (VSR) dataset Liu et al. (2022). Perception capability is tested on CIFAR-10/100 Krizhevsky et al. (2009) and MNIST LeCun (1998), following the guidance of Zhai et al. (2023). The SNLI-VE dataset Xie et al. (2019) evaluates the Visual Entailment capabilities, and POPE Li et al. (2023c) examines the tendency to objects hallucination.
@@ 6 Results and Discussion
Moreover, we visualize the task interference scores using Equation 3.2 and 3. Given that MixLoRA dynamically selects a subset of factors ($r$ out of $E$) for different instances, we record gradients concerning all $E$ factors and compare the task interference scores between standard LoRA models (with $r$=16) and MixLoRA (with $E$ = 16 and $r$ = 4). Figure 7 visualizes the interference scores for both LoRA $A$ and LoRA $B$ aggregated across all adaptation layers, including MLP and self-attention layers. The analysis reveals that MixLoRA ($E$=16, $r$=4) exhibits lower negative interference scores compared to the standard LoRA ($r$=16), underscoring MixLoRA’s efficacy in reducing task interference.